OpenStudy (anonymous):

Find the simple interest. principle $2000 Rate 5 1/2....months 9 what is the interest rounded to nearest cent??? please help

OpenStudy (anonymous):

I=P x R x T. 2000 x 5.5 x 9/12. Simplify. 2000 x 5.5 x 3/4=8250

OpenStudy (anonymous):

82.50?

OpenStudy (anonymous):

ty so much

OpenStudy (anonymous):

what formula are you using

OpenStudy (anonymous):

im cinfused I dont understand this so I dont know what formula to use

OpenStudy (anonymous):

we cannot direcxtly multiply percentage to a whole number, therefor let's convert percentage to decimal. 5.5% will be 5.5/100 my bad for multiplying 9 directly to the formula. here is the right way. for the t( number of years) , 9 months is 3/4 of a year. therefor the interest will be interest = 2000 ( 5.5/100) (3/4) = 82.5